A bag contains 20 tickets numbered from 1 to 20.Two tickets are drawn at random.The probability that both the numbers appearing on the tickets drawn will be prime:

Grade:12th Pass

1 Answers

ASh
6 years ago
Prime numbers between 1-20 = { 1, 3, 5, 7, 11, 13, 17, 19} 
Total number of Prime numbers = 8 
Total number of outlets = 20 
Therefore, 
P ( Prime number) = 8/20 
= 2/5 

Hope it helps you :)